OpenStudy (anonymous):

A certain drug is made from only two ingredients: compound A and compound B. There are 5 milliliters of compound A used for every 8 milliliters of compound B. If a chemist wants to make 1196 milliliters of the drug, how many milliliters of compound A are needed?

OpenStudy (alekos):

If we had 13 ml of the drug then there'd be 5 ml of A and 8 ml of B. So A is present in the ratio of 5:13 from here its just arithmetic

OpenStudy (shinalcantara):

say 'x' is the component used as a factor for both compounds. Since the total amount of mixture will be1196mL then 5x + 8x = 1196 Solve for the value of x After getting its value, you'll have compound A as 5x = 5 times the value of x

OpenStudy (hockeychick23):

5/8x+ 8/8x--> 13/8x= 1196 8/13(13/8x=1196) x=736 5/8(736)= 460 460+736= 1196 460 mL compound A 736 mL compound B